OpenStudy (anonymous):

What is the solution of 5a + 12 < 2 and 5a - 12 < 3?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

solve 5a + 12 < 2 for 'a' first

OpenStudy (anonymous):

I got a<-2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

solve 5a - 12 < 3 for 'a' next

OpenStudy (anonymous):

is that the whole answer:)

jimthompson5910 (jim_thompson5910):

no you need to solve 5a - 12 < 3 for 'a' now

OpenStudy (anonymous):

This is what it look like and I can only choose one answer: What is the solution of 5a + 12 < 2 and 5a - 12 < 3? A. a<3 B. a<2 C. a<-2\ D. a>-2 and btw for the second one I got A<3

jimthompson5910 (jim_thompson5910):

so the inequalities solve to a < -2 and a < 3 combine the two to get a < -2

jimthompson5910 (jim_thompson5910):

so it's C

OpenStudy (anonymous):

so i bring thos two answers together like this: a<-2 and a<3 and I get my answer?

jimthompson5910 (jim_thompson5910):

a < -2 and a < 3 combine to a < -2 think about it: if a is some number less than -2 AND less than 3, then it's just less than -2

OpenStudy (anonymous):

oh alright so my whole answer is c then right?

jimthompson5910 (jim_thompson5910):

yep after all that, the a < 3 is basically ignored example: say a = 2 sure 2 is less than 3 so that's why a = 2 works for a < 3 but a = 2 does NOT work for a < -2 so that's why a < -2 and a < 3 combine to a < -2

OpenStudy (anonymous):

oh okay thts wht I thought:)

OpenStudy (anonymous):

what is the solution of 1/2x-8=3/2x+2? I got x=5 is that right?

jimthompson5910 (jim_thompson5910):

can you draw it out? I'm not sure if the 1 is over the 2, 1 is over the 2x, or if 1is over the 2x-8

OpenStudy (anonymous):